OpenStudy (anonymous):

The degree measure of three angles of a triangle are x,y and z. If z= x+y/2 , hen find the value of z....pls HELP..

OpenStudy (anonymous):

Do you know the sum of all the angles of a triangle is how much??

OpenStudy (anonymous):

180

OpenStudy (anonymous):

Yes so you have: x + y + z = 180 Right??

OpenStudy (anonymous):

ya but wat does "degree measure" mean???

OpenStudy (anonymous):

So find z from here: z = 180 - (x + y) Also you are given with: \[z = \frac{x + y}{2}\] So: \[\large \frac{x + y}{2} = 180 - (x+y)\]

OpenStudy (anonymous):

Now find (x+y) Multiply 2 both the sides: \[\large x+ y = 360 - 2(x + y)\] \[\large (x+y) + 2(x + y) = 360\] \[\large 3(x+y) = 360\] Divide by 3 both the sides: \[\large x + y = 120\]

OpenStudy (anonymous):

So: \[\large z = \frac{x+y}{2} = \frac{120}{2} = ??\]

OpenStudy (anonymous):

60.....thnx

OpenStudy (anonymous):

Welcome dear...